OpenStudy (anonymous):

I need help with this question... I understand that 4^-2 is 1/4^2, but I don't know how I would simplify the rest of the equation. Thank you in advance!!! 4^-2 x 8^2/3 x 4^3/2

OpenStudy (anonymous):

ok the exponent 2/3 means ' the cube root of 8 squared'. with fractional exponents the top is the power and the bottom is a root. so 8^2/3 = 2 (the cube root of 8) ^2 = 2^2 = 4 can you figure out 4^ 3/2?

OpenStudy (anonymous):

4^ 3/2 = (sqrt 4) cubed

OpenStudy (anonymous):

do you follow that ok?

OpenStudy (anonymous):

Wow you are incredible! Yes, I understand that perfectly. So it would cube the whole square root? The answer would be 8 because the square root of 4 is 2, then cube it is 8.

OpenStudy (anonymous):

dead right!